Projection Close 11/7



To do this evening/ was completed: No issues this evening, but I forgot to mention the other night that I noticed a hole in the ceiling in the lobby (that the electricians had left exposed with its cover off to the side), so I pulled it over to cover it up God forbid anything were to fall from the ceiling and hit a patron... 

Presentation Issues:

Preshow Clips/Playlist: Built the A1 playlists for the upcoming week of 11/13. We're still waiting on some content. 

Trailer Updates:

Special Events/ AV set up: Continued DCPing NoBudge shorts, DCP'd three of them this evening. They're all transcoding to the NoBudge folder in the Share folder. There is only one more to do, which is Plastic Glass. 

We received a BluRay of Mad Max for a private event on 11/21. Ripped it in T1 and ran it through Handbrake. We can DCP this tomorrow. 

Ordering/Receiving (posters & trailers):

Ordering/Receiving (content):

Outgoing/Shipping (content):

Print/video Inspection: Started inspecting The Lord of the Rings. Report is still up in T1, so please feel free to add to that tomorrow. 

Asset tracking doc updated/checked?:

Maintenance/cleaning:

Cue sheets updated/checked (every shift): Yes

Marquee updated?:

Posters updated?:

Cleaning lights on for the night?:

Projection cart by entrance?:

Additional Notes:

Comments

Popular posts from this blog

Projection Admin 10/2

Projection Admin 9/18

Projection Admin 9/2